How much matter needs to be converted to energy to supply 400 kJ of energy, the approximate energy of 1 mol of C–H bonds? What conclusion does this suggest about energy changes of chemical reactions?
Answer #1

Solution:-

E = mc2 .........................(1)

Where , E = energy

m = mass

c = speed of light (= 3 × 108 m/ s)

E = 400 kJ = 4 × 105 J (given)

Putting values of E and c , in equation (1), we get

4 × 105 = m × (3 × 108)2

or, m = (4 × 105) / (9 × 1016) = 4.44 × 10-12 kg

Hence, 4.44 × 10-12 kg of matter required to be converted to energy to supply 400 kJ of energy.

--------------------------------------------------------------